# Movable Antennas Meet Intelligent Reflecting Surface: When Do We Need Movable Antennas?

### v1

@@ Abstract
Intelligent reflecting surface (IRS) and movable antenna (MA)/fluid antenna (FA) techniques have both received increasing attention in the realm of wireless communications due to their ability to reconfigure and improve wireless channel conditions. In this paper, we investigate the integration of MAs/FAs into an IRS-assisted wireless communication system. In particular, we consider the downlink transmission from a multi-MA base station (BS) to a single-antenna user with the aid of an IRS, aiming to maximize the user’s received signal-to-noise ratio (SNR), by jointly optimizing the BS/IRS active/passive beamforming and the MAs’ positions. Due to the similar capability of MAs and IRS for channel reconfiguration, we first conduct theoretical analysis of the performance gain of MAs over conventional fixed-position antennas (FPAs) under the line-of-sight (LoS) BS-IRS channel and derive the conditions under which the performance gain becomes more or less significant. Next, to solve the received SNR maximization problem, we propose an alternating optimization (AO) algorithm that decomposes it into two subproblems and solve them alternately. Numerical results are provided to validate our analytical results and evaluate the performance gains of MAs over FPAs under different setups.
@@ I Introduction
The intelligent reflecting surface (IRS) is regarded as a promising technique for future sixth-generation (6G) wireless networks due to its superior characteristics such as channel reconfiguration capability, low-power consumption and low-cost deployment [1]. By individually adjusting the reflecting coefficient of each reflecting element, the IRS can jointly alter the strength/direction of its reflected signal for achieving various objectives, such as signal enhancement, interference nulling, spatial multiplexing among them [2, 3].
On the other hand, the movable antenna (MA)/fluid antenna (FA) technology has drawn great interest in academia and industry recently. Similar to the IRS, the MA/FA technology can reconfigure wireless channel conditions by enabling multiple antennas to be flexibly moved within a confined region at the transmitter/receiver [4, 5, 6]. As such, both IRS and MA/FA technologies have been extensively studied in the literature under various system setups, e.g., multiple-input single-output (MISO) [7, 8, 9], multiple-input multiple-output (MIMO) [10, 11], spectrum sharing [12, 13], physical-layer security [14, 15], and so on. Furthermore, some recent works have also proposed a new MA-inspired IRS architecture with movable reflecting elements and characterized its performance limits [16, 17]. However, to the best of our knowledge, there is no existing works focusing on the integration of MAs into an IRS-assisted wireless system. Particularly, due to the comparable capability of MAs and IRS for channel reconfiguration, it remains an open problem whether the performance gain of MAs still exists or not in the presence of the IRS’s passive beamforming.
@@
To fill in this gap, we investigate in this paper an MA-enhanced IRS-assisted MISO communication system, where a base station (BS) equipped with multiple MAs communicates with a user with the aid of a fixed IRS. We aim to maximize the received signal-to-noise ratio (SNR) at the user by jointly optimizing the active/passive beamforming at the BS/IRS and the MAs’ positions, which appears to be a highly non-convex problem that is challenging to be optimally solved. Due to the similar capabilities of MAs/IRS for channel reconfiguration, we first conduct theoretical analysis of the performance gain of MAs over conventional fixed-position antennas (FPAs) under the line-of-sight (LoS) BS-IRS channel to drive essential insights. It is unveiled that the performance gain of MAs over FPAs vanishes in the case of a single MA or far-field BS-IRS channel condition. Next, we propose an alternating optimization (AO) algorithm to solve the primal SNR maximization problem by decomposing it into two subproblems and solve them alternately. Numerical results are provided to validate our theoretical analysis and evaluate the performance gains of MAs over FPAs under different setups.
@@ II System Model and Problem Formulation
$||\boldsymbol{t}_{i}-\boldsymbol{t}_{j}||_{2}\geq D_{\min},\forall i,j\in\mathcal{N},i\neq j,$ (3c)
@@
where $D_{\min}$ denotes the minimum spacing between any two MAs to avoid mutual antenna coupling. Here, we assume that all required channel state information are available by applying the existing channel estimation techniques dedicated to MAs/IRSs, e.g., [19] and [20].
However, (P1) is a non-convex optimization problem that is challenging to solve due to the spacing constraints for MAs (i.e., (3c)) and the unit-modulus constraints for IRS passive beamforming (i.e., (3a)). To reveal essential insights, we first conduct performance analysis in the next section to evaluate the fundamental gains of integrating MAs into an IRS-aided wireless communication system.
@@ III Performance Analysis: MAs versus FPAs
To facilitate our performance analysis, we first consider that the BS is equipped with a single MA, i.e., $N=1$, and characterize the maximum BS-user end-to-end channel power gain over different positions within the transmit region $\mathcal{C}_{t}$ under the optimal IRS passive beamforming. As a result, the APV reduces to a column vector, $\boldsymbol{t}\in\mathbb{R}^{3\times 1}=[x_{t},y_{t},z_{t}]^{T}$. Moreover, we consider that the IRS can achieve an LoS-dominant channel with the BS and be located in its near-field region. Moreover, we consider that the IRS can achieve an LoS-dominant channel with the BS, which usually hold in practice if the IRS is deployed in the vicinity of the BS [18]. To capture the close BS-IRS distance in this case, we consider in this section a general near-field BS-IRS channel model, which is given by [21]
@@ III-A Optimal Antenna Position
To verify the above analysis, we plot in Fig. 2(a) the received SNR at the user versus the BS-IRS distance under the near-field LoS BS-IRS channel. The BS is equipped with a single-MA. The operating frequency is $f=5$ GHz. The total number of the IRS reflecting elements is $M=25^{2}$. The transmit SNR is $P/\sigma^{2}=110$ dB. In the FPA benchmark scheme, the antenna is fixed at (10). It is observed that the SNR performance by a single MA and a single FPA are identical for all BS-IRS distances considered, which validates our analysis.
@@
where $\boldsymbol{q}_{U}=[x_{U},y_{U},z_{U}]^{T}\in\mathbb{R}^{3}$ denotes the coordinate of the user. It can be verified that the optimal antenna position that maximizes (11) is given by
@@ III-B Fluctuation of Channel Power Gain
Based on (14), we analyze the effects of different key system parameters on it. First, it is easy to see that $H(M_{y},M_{z},\boldsymbol{t}_{1})\pm H(M_{y},M_{z},\boldsymbol{t}_{2})$ monotonically increases with $M_{y}$ and $M_{z}$, provided that $R(\boldsymbol{t}_{1})<R(\boldsymbol{t}_{2})$. Thus, $G_{u}(\boldsymbol{t}_{1})-G_{u}(\boldsymbol{t}_{2})$ also monotonically increases with $M_{y}$ and $M_{z}$, for any given $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$. This indicates that by increasing the size of the IRS, the maximum channel power gains within $\mathcal{C}_{t}$ experiences more significant fluctuation.
@@
Based on the above, although the antenna position yielding the maximum channel power gain is fixed as (10), the channel power gain within $\mathcal{C}_{t}$ may vary. As such, it can be inferred that in the case of multiple antennas, the performance gain of MAs over FPAs may still exist, unlike the single-MA case. To verify this claim, we plot in Fig. 2(b) the received SNR at the user versus the BS-IRS distance with $N=4$ MAs, with other simulation parameters the same as Fig. 2(a). It is observed that different from the single-MA case, employing multiple MAs can still yield a performance gain over FPAs, if the BS-IRS distance is small. Moreover, the performance gain is observed to decrease with the BS-IRS distance, which is consistent with our previous analysis.
@@ IV Proposed Solution to (P1)
$\boldsymbol{w}(\boldsymbol{T},\boldsymbol{\Psi})=\sqrt{P}\frac{\left(\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Psi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\right)^{H}}{||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Psi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||}.$ (17)
@@ IV-A Optimizing $\boldsymbol{\Psi}$ for Given $\boldsymbol{T}$
where $\boldsymbol{\alpha}_{m}=\sum_{i\neq m}^{M}{\boldsymbol{g}_{1,i}e^{-j\varphi_{i}}}$. As such, in the $m$-th BCD iteration, we can fix $\varphi_{m}$, $m\neq i$, $m\in{\cal M}$ and optimize $\varphi_{m}$ as (20). Then, the BCD iteration proceeds until the phase shifts of all $M$ IRS reflecting elements have been updated.
@@ IV-B Optimizing $\boldsymbol{T}$ for Given $\boldsymbol{\Psi}$
which can be solved optimally by applying the graph-based approach proposed in our previous work [9] with slight modification and modifying the weight assignment therein.
@@ V Numerical Results
In this section, we provide numerical results to evaluate the efficacy of our proposed AO algorithm. Unless other specified, the simulation parameters are set as follows. The operating frequency is $f=5$ GHz. The number of the transmit MAs is $N=4$. The total number of the IRS reflecting elements along the $y$- and $z$-axes is set identical as $M_{y}=M_{y}=15$, leading to $M=M_{y}\times M_{z}=225$. The spacing between any two adjacent IRS reflecting elements is $d=\frac{\lambda}{2}$, and the minimum spacing between any two MAs at the BS is $D_{\min}=\frac{\lambda}{2}$. The length of the transmit region $\mathcal{C}_{t}$ is $A=0.6$ m. The spacing of the sampling points is $\delta_{s}=0.01$ m. The transmit SNR is $P/\sigma^{2}=110$ dB, and the coordinate of the center of the transmit region $\mathcal{C}_{t}$ is $\boldsymbol{q}_{B}=[5,5,0]^{T}$. We consider the Rician-fading IRS-user channel with the Rician factor set to $K=3$ dB. Ths distance from the IRS to the user is $d_{\text{IU}}=30$ m, and the path loss exponent for the IRS-user channel is $\kappa=2.8$. Moreover, we consider the near-field multi-path BS-IRS channel. Let $L$ denote the number of dominant paths over the BS-IRS link, and $\boldsymbol{s}_{\text{BI},l}\in\mathbb{R}^{3\times 1}$ denote the coordinate of the $l$-th scatterer, $l=1,2,\cdots,L$. Then, the BS-IRS channel can be expressed as [21]
@@
Figure 3: Received SNR at the user versus BS-IRS distance.
@@
Next, we plot in Fig. 4 the received SNR at the user versus the number of the transmit MAs (i.e., $N$) with $L=8$. It is observed that the received SNRs by all schemes increase with the number of MAs, thanks to the enlarged beamforming gain. Moreover, as $N$ increases, employing MAs is observed to yield more significant gain over the two benchmarks, implying that more spatial diversity gain can be exploited for a larger $N$ in the case of the multi-path BS-IRS channels.
Lastly, we plot in Fig. 5 the received SNR versus the length of the transmit MA array, i.e., $A$, with $N=4$ and $L=8$. It is observed that the received SNRs by all schemes increase with $M$, as more reflecting elements can provide more significant passive beamforming gain. However, the performance gain of the proposed AO algorithm over the FPA w/o AS slightly decreases with $M$, e.g., from 2.01 dB at $M=9^{2}$ to 1.58 dB at $M=19^{2}$. This observation is also similar to the trend for the BS-IRS LoS channel discussed in Section III-B.
@@ VI Conclusions
In this paper, we studied a joint active/passive beamforming and MA position optimization problem for an MA-enhanced IRS-assisted wireless communication system and proposed an AO algorithm to solve it. Our performance analysis shows that under the LoS BS-IRS channel and optimal IRS passive beamforming, the performance gain of MAs over FPAs vanishes in the case of a single MA or far-field BS-IRS propagation, but increases as the BS-IRS channel approaches the near-field. Numerical results also revealed a similar trend for multi-path BS-IRS channels and that MAs achieve more significant gains over FPAs compared to LoS BS-IRS channels.

### v3

@@ Abstract
Intelligent reflecting surface (IRS) and movable antenna (MA)/fluid antenna (FA) techniques have both received increasing attention in the realm of wireless communications due to their ability to reconfigure and improve wireless channel conditions. In this paper, we investigate the integration of MAs/FAs into an IRS-assisted wireless communication system. In particular, we consider the downlink transmission from a multi-MA base station (BS) to a single-antenna user with the aid of an IRS, aiming to maximize the user’s received signal-to-noise ratio (SNR), by jointly optimizing the BS/IRS active/passive beamforming and the MAs’ positions. Due to the similar capability of MAs and IRS for channel reconfiguration, we first conduct theoretical analyses of the performance gain of MAs over conventional fixed-position antennas (FPAs) under the line-of-sight (LoS) BS-IRS channel and derive the conditions under which the performance gain becomes more or less significant. Next, to solve the received SNR maximization problem, we propose an alternating optimization (AO) algorithm that decomposes it into two subproblems and solve them alternately. Numerical results are provided to validate our analytical results and evaluate the performance gains of MAs over FPAs under different setups.
@@ I Introduction
The intelligent reflecting surface (IRS) is regarded as a promising technique for future sixth-generation (6G) wireless networks due to its superior characteristics such as channel reconfiguration capability, low-power consumption, and low-cost deployment [1]. By individually adjusting the reflecting coefficient of each reflecting element, the IRS can jointly alter the strength/direction of its reflected signal for achieving various objectives, such as signal enhancement, interference nulling, and spatial multiplexing among them [2, 3].
On the other hand, the movable antenna (MA)/fluid antenna (FA) technology has drawn great interest in academia and industry recently. Similar to the IRS, the MA/FA technology can reconfigure wireless channel conditions by enabling multiple antennas to be flexibly moved within a confined region at the transmitter/receiver [4, 5, 6]. As such, both IRS and MA/FA technologies have been extensively studied in the literature under various system setups, e.g., multiple-input single-output (MISO) [7, 8, 9], multiple-input multiple-output (MIMO) [10, 11], spectrum sharing [12, 13], physical-layer security [14, 15], and so on. Furthermore, some recent works have also proposed a new MA-inspired IRS architecture with movable reflecting elements and characterized its performance limits [16, 17]. However, to the best of our knowledge, there are no existing works focusing on the integration of MAs into an IRS-assisted wireless system. Particularly, due to the comparable capability of MAs and IRS for channel reconfiguration, it remains an open problem whether the performance gain of MAs still exists or not in the presence of the IRS’s passive beamforming.
@@
To fill in this gap, we investigate in this paper an MA-enhanced IRS-assisted MISO communication system, where a base station (BS) equipped with multiple MAs communicates with a user with the aid of a fixed IRS. We aim to maximize the received signal-to-noise ratio (SNR) at the user by jointly optimizing the active/passive beamforming at the BS/IRS and the MAs’ positions, which appears to be a highly non-convex problem that is challenging to be optimally solved. Due to the similar capabilities of MAs/IRS for channel reconfiguration, we first conduct theoretical analyses of the performance gain of MAs over conventional fixed-position antennas (FPAs) under the line-of-sight (LoS) BS-IRS channel to drive essential insights. It is unveiled that the performance gain of MAs over FPAs vanishes in the case of a single MA or far-field BS-IRS channel condition. Next, we propose an alternating optimization (AO) algorithm to solve the primal SNR maximization problem by decomposing it into two subproblems and solve them alternately. Numerical results are provided to validate our theoretical analyses and evaluate the performance gains of MAs over FPAs under different setups.
@@ II System Model and Problem Formulation
$||\boldsymbol{t}_{i}-\boldsymbol{t}_{j}||\geq D_{\min},\forall i,j\in\mathcal{N},i\neq j,$ (3c)
@@
where $D_{\min}$ denotes the minimum spacing between any two MAs to avoid mutual antenna coupling. Here, we assume that all required channel state information is available by applying the existing channel estimation techniques dedicated to MAs/IRSs, e.g., [18] and [19].
However, (P1) is a non-convex optimization problem that is challenging to solve due to the spacing constraints for MAs (i.e., (3c)) and the unit-modulus constraints for IRS passive beamforming (i.e., (3a)). To reveal essential insights, we first conduct performance analyses in the next section to evaluate the fundamental gains of integrating MAs into an IRS-aided wireless communication system.
@@ III Performance Analysis: MAs versus FPAs
To facilitate our performance analyses, we first consider that the BS is equipped with a single MA, i.e., $N=1$, and characterize the maximum BS-user end-to-end channel power gain over different positions within the transmit region $\mathcal{C}_{t}$ under the optimal IRS passive beamforming. As a result, the APV reduces to a column vector, $\boldsymbol{t}\in\mathbb{R}^{3\times 1}=[x_{t},y_{t},z_{t}]^{T}$. Moreover, we consider that the IRS can achieve a LoS-dominant channel with the BS, which usually holds in practice if the IRS is deployed in the vicinity of the BS [20]. To capture the close BS-IRS distance in this case, we consider in this section a general near-field BS-IRS channel model, which is given by [21]
@@ III-A Optimal Antenna Position
To verify the above analyses, we plot in Fig. 2(a) the received SNR at the user versus the BS-IRS distance under the near-field LoS BS-IRS channel. The BS is equipped with a single-MA. The operating frequency is $f=5$ GHz. The total number of the IRS reflecting elements is $M=25^{2}$. The transmit SNR is $P/\sigma^{2}=110$ dB. In the FPA benchmark scheme, the antenna is fixed at (10). It is observed that the SNR performance by a single MA and a single FPA are identical for all BS-IRS distances considered, which validates our analyses.
@@
where $\boldsymbol{q}_{U}=[x_{U},y_{U},z_{U}]^{T}\in\mathbb{R}^{3\times 1}$ denotes the coordinate of the user. It can be verified that the optimal antenna position that maximizes (11) is given by
@@ III-B Fluctuation of Channel Power Gain
Based on (14), we analyze the effects of different key system parameters on it. First, it is easy to see that $H(M_{y},M_{z},\boldsymbol{t}_{1})\pm H(M_{y},M_{z},\boldsymbol{t}_{2})$ monotonically increases with $M_{y}$ and $M_{z}$, provided that $R(\boldsymbol{t}_{1})<R(\boldsymbol{t}_{2})$. Thus, $G_{u}(\boldsymbol{t}_{1})-G_{u}(\boldsymbol{t}_{2})$ also monotonically increases with $M_{y}$ and $M_{z}$, for any given $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$. This indicates that by increasing the size of the IRS, the maximum channel power gains within $\mathcal{C}_{t}$ experience more significant fluctuation.
@@
Based on the above, although the antenna position yielding the maximum channel power gain is fixed as (10), the channel power gain within $\mathcal{C}_{t}$ may vary. As such, it can be inferred that in the case of multiple antennas, the performance gain of MAs over FPAs may still exist, unlike the single-MA case. To verify this claim, we plot in Fig. 2(b) the received SNR at the user versus the BS-IRS distance with $N=4$ MAs and other simulation parameters the same as Fig. 2(a). It is observed that different from the single-MA case, employing multiple MAs can still yield a performance gain over FPAs, if the BS-IRS distance is small. Moreover, the performance gain is observed to decrease with the BS-IRS distance, which is consistent with our previous analyses.
@@ IV Proposed Solution to (P1)
$\boldsymbol{w}(\boldsymbol{T},\boldsymbol{\Psi})=\frac{\left(\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Psi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})\right)^{H}}{||\boldsymbol{h}_{\text{IU}}^{H}\boldsymbol{\Psi}\boldsymbol{H}_{\text{BI}}(\boldsymbol{T})||}.$ (17)
@@ IV-A Optimizing $\boldsymbol{\Psi}$ for Given $\boldsymbol{T}$
where $\boldsymbol{\alpha}_{m}=\sum_{i\neq m}^{M}{\boldsymbol{g}_{1,i}e^{-j\varphi_{i}}}$. As such, in the $m$-th BCD iteration, we can fix $\varphi_{i}$, $i\neq m$, $i\in{\cal M}$ and optimize $\varphi_{m}$ as (20). Then, the BCD iteration proceeds until the phase shifts of all $M$ IRS reflecting elements have been updated.
@@ IV-B Optimizing $\boldsymbol{T}$ for Given $\boldsymbol{\Psi}$
which can be solved optimally by applying the graph-based approach proposed in our previous work [9] with modifying the weight assignment therein.
@@ V Numerical Results
In this section, we provide numerical results to evaluate the efficacy of our proposed AO algorithm. Unless other specified, the simulation parameters are set as follows. The operating frequency is $f=5$ GHz. The number of the transmit MAs is $N=4$. The total number of the IRS reflecting elements along the $y$- and $z$-axes is set identical as $M_{y}=M_{y}=15$, leading to $M=M_{y}\times M_{z}=225$. The spacing between any two adjacent IRS reflecting elements is $d=\frac{\lambda}{2}$, and the minimum spacing between any two MAs at the BS is $D_{\min}=\frac{\lambda}{2}$. The length of the transmit region $\mathcal{C}_{t}$ is $A=0.6$ m. The spacing of the sampling points is $\delta_{s}=0.01$ m. The transmit SNR is $P/\sigma^{2}=110$ dB, and the coordinate of the center of the transmit region $\mathcal{C}_{t}$ is $\boldsymbol{q}_{B}=[5,5,0]^{T}$. We consider the Rician-fading IRS-user channel with the Rician factor set to $K=3$ dB. The distance from the IRS to the user is $d_{\text{IU}}=30$ m, and the path loss exponent for the IRS-user channel is $\kappa=2.8$. Moreover, we consider the near-field multi-path BS-IRS channel. Let $L$ denote the number of dominant paths over the BS-IRS link, and $\boldsymbol{s}_{\text{BI},l}\in\mathbb{R}^{3\times 1}$ denote the coordinate of the $l$-th scatterer, $l=1,2,\cdots,L$. Then, the BS-IRS channel can be expressed as [21]
@@
Figure 3: Received SNR at the user versus the BS-IRS distance.
@@
Next, we plot in Fig. 4 the received SNR at the user versus the number of the transmit MAs (i.e., $N$) with $L=8$. It is observed that the received SNRs by all schemes increase with the number of MAs, thanks to the enlarged beamforming gain. Moreover, as $N$ increases, employing MAs is observed to yield a more significant gain over the two benchmarks, implying that more spatial diversity gain can be exploited for a larger $N$ in the case of the multi-path BS-IRS channels.
Lastly, we plot in Fig. 5 the received SNR versus the number of IRS reflecting elements, i.e., $M$, with $N=4$ and $L=8$. It is observed that the received SNRs by all schemes increase with $M$, as more reflecting elements can provide more significant passive beamforming gain. However, the performance gain of the proposed AO algorithm over the FPA w/o AS slightly decreases with $M$, e.g., from 2.01 dB at $M=9^{2}$ to 1.58 dB at $M=19^{2}$. This observation is also similar to the trend for the BS-IRS LoS channel discussed in Section III-B.
@@ VI Conclusions
In this paper, we studied a joint active/passive beamforming and MA position optimization problem for an MA-enhanced IRS-assisted wireless communication system and proposed an AO algorithm to solve it. Our performance analyses shows that under the LoS BS-IRS channel and optimal IRS passive beamforming, the performance gain of MAs over FPAs vanishes in the case of a single MA or far-field BS-IRS propagation, but increases as the BS-IRS channel approaches the near-field. Numerical results also revealed a similar trend for multi-path BS-IRS channels and that MAs achieve more significant gains over FPAs compared to LoS BS-IRS channels.